OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places. 11 pi/ 4

OpenStudy (anonymous):

@freckles ?

OpenStudy (anonymous):

multiply by 180, divide by \(\pi\) i.e.. erase the \(\pi\) and multiply by 180

OpenStudy (anonymous):

495? why did you mult. 180?

OpenStudy (anonymous):

\[\frac{(11 \pi ) 180}{4 \pi }=495 {}^{\circ} \]

OpenStudy (anonymous):

pi =180 degrees Dividing by pi is finding out how many pi s are in the original expression.